# Deep learning inference of the neutron star equation of state

### v1

@@ I Introduction
Recently, machine learning techniques have also been used to predict EOS’ features from observations [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Roughly, two different approaches have been followed in the literature: reconstructing the EOS from either stellar masses, stellar radii and tidal deformabilities [21, 22, 24, 25, 26, 28, 29, 30, 31, 33, 35, 37, 39, 40, 41, 43, 44, 45], or focusing directly on the nuclear matter saturation properties [23, 27, 32, 34, 36, 38, 41, 42]. Our work falls into the first category. In particular, we construct a pipeline for supervised deep neural networks which allow us to predict the EOS from mock observational data composed of masses, radii and tidal deformabilities (or equivalently, tidal Love number). We employ both conventional (deterministic) and Bayesian (probabilistic) architectures. The latter utilises variational layers, where the weights and biases are drawn from a learnable posterior distribution, whereas in the deterministic network the weights and biases are point estimates. To the best of our knowledge, the employment of variational layers to account for epistemic uncertainty in the dataset is a novelty. While Bayesian neural networks have previously been used in the literature, our approach is inherently different. In particular, in Refs. [41, 42], the authors employed a Bayesian approach to minimise a log-likelihood related to neutron star observables, using a fundamentally different EOS parametrization to ours. What is more, Ref. [45] studied inferences of the EOS directly from telescope spectra employing the method of normalising flow for the reconstruction of the probability density functional.
We work in the context of a new prescriptions for modelling EOSs, following Ref. [46]. We assume that neutron stars are described by an EOS composed of three regions. At low-density we employ a realistic tabulated EOS (AP4 or SLy), while at high-density we introduce a QCD phase by means of a speed-of-sound parametrization [47]. Most importantly, we also allow for a vacuum energy phase transition in the stellar core. The latter is a key novelty of our work. A vacuum energy shift, which can be interpreted as a screening of the underlying cosmological constant, is to be expected in the core of the star, where the pressure can be sufficiently high to trigger the QCD phase transition. This mechanism has been proposed in Ref. [46] to study neutron stars as novel probes for the cosmological constant problem.
@@
The paper is structured as follows: in Section II we describe the modelling of the neutron star interior and the numerical pipeline we use to solve the equations which leads to our training and testing datasets. In Section III we describe our deep-learning pipeline, which we split into an analysis based on a deterministic and a Bayesian deep network respectively. We summarise in Section IV. The numerical data and codes of our analysis are available online at Ref. [48].
Figure 1: A sketch showing an overview of the deep networks we use, i.e. a “standard” deterministic and a probabilistic network respectively. The input data to our network are simulated values for the mass, radius and tidal number of the given neutron star, computed according to the procedure explained in the paper, whereas the output is predominantly a prediction of the speed of sound. Our probabilistic deep network accounts for the epistemic uncertainty in the dataset, by using learnable probability functions for the weights and biases of the network, as explained in Sec. III.3.
II Neutron star modelling
@@
$\begin{split}m^{\prime}(r)&=4\pi r^{2}\epsilon(r)\ ,\\ p^{\prime}(r)&=-\frac{p(r)+\epsilon(r)}{r[r-2m(r)]}G[m(r)+4\pi r^{3}p(r)],\\ \nu^{\prime}(r)&=-\frac{2p^{\prime}(r)}{p(r)+\epsilon(r)}\ ,\end{split}$ (2)
@@
$Q_{ij}=-\lambda\mathcal{E}_{ij}=-\frac{2}{3}k_{2}R^{5}\mathcal{E}_{ij}\ ,$ (3)
where in the last step we have introduced the dimensionless $l=2$ Love number, $k_{2}$ [51, 52, 53, 54, 55]. The latter can be computed by considering a linear perturbation to the background solution (1),
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+h_{\mu\nu}(r,\theta,\phi)\ ,$ (4)
@@
$h_{\mu\nu}=$ $e^{\nu(r)}H(r)Y_{20}(\theta,\phi)dt^{2}+e^{\mu(r)}H(r)Y_{20}(\theta,\phi)dr^{2}$
$+r^{2}K(r)Y_{20}(\theta,\phi)d\Omega^{2}\ .$ (5)
$T_{\mu\nu}=T_{\mu\nu}^{(0)}+\delta T_{\mu\nu},$ (6)
@@
$H^{\prime\prime}(r)-\frac{2}{r}e^{\mu}\big{[}2\pi Gr^{2}(\epsilon-p)-1\big{]}H^{\prime}(r)\\ -2e^{\mu}\bigg{\{}\frac{3}{r^{2}}-2\pi G[5\epsilon+9p+(\epsilon+p)f]\\ +2G^{2}e^{\mu}\left(\frac{m(r)}{r^{2}}+4\pi rp\right)^{2}\bigg{\}}H(r)=0\ ,$ (7)
where $f=d\epsilon/dp$. Introducing $\beta(r)\equiv dH/dr$, we solve Eq. (7) integrating from the stellar center up to the surface with boundary conditions. In the inner boundary we impose
$H(r_{0})\sim a_{0}r_{0}^{2}\quad\ ,\quad\beta(r_{0})=2a_{0}r_{0}\ ,$ (8)
@@
$H(r\geq R)=c_{1}{\cal P}_{22}\left(r/M-1\right)+c_{2}{\cal Q}_{22}\left(r/M-1\right)\ ,$ (9)
@@
$-\frac{1+g_{tt}}{2}=-\frac{m}{r}-\frac{3}{2}\frac{Q_{ij}}{r^{3}}n^{i}n^{j}+...+\frac{\mathcal{E}_{ij}}{2}r^{2}n^{i}n^{j}+...,$ (10)
@@
$k_{2}=\frac{8C^{5}}{5}(1-2C)^{2}(2+2C(y-1)-1)\\ \times\bigg{\{}2C\big{[}6-3y+3C(5y-8)\big{]}\\ +4C^{3}\big{[}13-11y+C(3y-2)+2C^{2}(1+y)\big{]}\\ +3(1-2C)^{2}\big{[}2-y+2C(y-1)\big{]}\text{ln}(1-2C)\bigg{\}}^{-1},$ (11)
@@
In summary, given a choice for the EOS, solving the system in Eq. (2) together with Eq. (7) allows us to extract the total stellar radius $R$, mass $M$ and dimensionless $l=2$ Love number $k_{2}$ respectively. Schematically,
$(\text{EOS})\mapsto(M,\,R,\,k_{2}).$ (12)
@@
A high-density regime for $\rho>\rho_{\text{tr}}$, where the nuclear matter experiences a QCD phase transition. We employ a phenomenological EOS using an agnostic speed-of-sound model [63, 47]. In particular, we construct a linear approximation of the speed of sound $c_{s}^{2}\equiv\frac{\partial p(\epsilon)}{\partial\epsilon}$, randomly sampling seven reference points $(\rho,c_{s})$ within $\rho\in(\rho_{\text{tr}},12\,\rho_{0})$ and $c_{s}\in(0,1)$, and connecting them by linear segments. Note that for the seventh point, we always assume the mass density to be fixed, i.e. $\rho_{7}=12\,\rho_{0}$, while we allow the speed of sound to acquire random values. We assume causality is preserved, thus we discard models with $c_{s}>1$ to avoid unphysical EOSs.
@@
a vacuum energy phase transition in the inner stellar core, for values of the pressure larger than $p_{c}=(200\,\text{MeV})^{4}$. Indeed, deep within the stellar core, the local pressure can be sufficiently high to trigger the QCD phase transition. This would generate a shift in the underlying value of the vacuum energy. One can interpret this shift in energy as a new effective cosmological constant term $\Lambda$ for the new exotic phase. We follow Ref. [46], and we choose 10 values of $\Lambda$ in the range $\big{(}-(194\,\text{MeV})^{4},(194\,\text{MeV})^{4}\big{)}$, including the case where no phase transition develops in the core, i.e. $\Lambda=0$. We allow for both positive and negative values of $\Lambda$, to consider the possibility of an unknown particle physics mechanism screening some or all of the underlying change in vacuum energy.
III Deep learning pipeline
@@
In this section, we describe the concrete setup of our deep learning network. Our scope is to constrain the neutron star EOS from observable properties, such as the mass, the radius, and the tidal deformability of the star, or equivalently its $l=2$ dimensionless tidal Love number. Doing so corresponds to solving the “inverse” problem we delineated in Sec. II and described by Eq. (12), that is, we aim at computing
$(M,\,R,\,k_{2})\mapsto(\text{EOS}).$ (13)
@@
While it is known that such inverted mapping exists [64], one cannot practically retrieve it. First of all, we cannot form a continuous curve from the $M$-$R$ observational data or the $k_{2}$-$M$ ones. Furthermore, observational data have uncertainties and are described by distributions, hence one is not able to reconstruct the genuine $M$-$R$ and $k_{2}$-$M$ relations (see Ref. [26] for a further discussion on the issue). Lastly, some initial information about the EOS at low density should be known beforehand. We tackle this challenge with an alternative approach, similar to Ref. [26]. In particular, instead of seeking the unique inverse mapping, i.e. Eq. (13), we infer the most likely EOS from observations data including their uncertainties through a regression analysis with deep-learning techniques. We do this at two steps. We first consider a “standard” (deterministic) deep network. In this case, the weights and biases of the network at training are point estimates, which implies that once the network is trained, a given input will always output the same prediction, modulo the inherit stochasticity in deep networks. Our second approach will be a Bayesian deep network, able to capture the epistemic uncertainty in the data. In this approach, the weights and biases at each node are drawn from learnable probability density functions, and the output of the network will be a statistical realisation of the given input. Before we dive into the description of the deep architecture and their results, we first discuss the construction of the training and validation dataset.
III.1 Training and validation dataset
@@
In order to train and test both the deterministic and the Bayesian neural networks, we first generate mock observational data as input data. We first solve the TOV equations (2) for randomly generated EOSs (see Sec. II), obtaining for each EOS a $M$-$R$ and a $k_{2}$-$M$ curve respectively. Note that, in order to generate data in agreement with the heaviest neutron star observed so far (PSR J0952-0607 [65]), which has a mass $M_{\text{max}}=(2.35\pm 0.17)M_{\odot}$, we discard all EOSs which predict a maximum mass not included in this interval. For each EOS, we then randomly sample $N$ mock observation points $(M_{i},R_{i},k_{2,i})$, assuming a uniform distribution of neutron star masses in the interval $(M_{\text{min}},M_{\text{max}})$, where $M_{\text{min}}$ is the minimum mass produced by the given EOS  [^2] The choice of this interval allow us to discard unstable branches of the $M$-$R$ curve, while associating each $M$ to a single value of $R$ and $k_{2}$.. We initially chose $N=15$, following Ref. [26], but we also took into consideration the case of a larger number of mock data for a given EOS, i.e. we consider $N=20$ and $N=30$. As we will discuss in Sec. III.2, allowing for a larger set of observation points for each $M$-$R$ and $k_{2}$-$M$ curve increases the accuracy of our network.
@@
As we discussed in Sec. II, we generate $2\times 10^{4}$ EOSs, of which only $1491$ survive the maximum mass test. For each set of points $N\times(M^{\ast}_{i},R^{\ast}_{i},k^{\ast}_{2,i})$, we create $n$ different realisations through noise injection, leading to $N\times n\times 3$ points in total for each EOS. All values are normalized appropriately by employing $M^{\ast}_{i}/M_{\text{norm}}$, and $R^{\ast}_{i}/R_{\text{norm}}$ with $M_{\text{norm}}=3\,M_{\odot}$ and $R_{\text{norm}}=20\,\text{km}$ for masses and radii respectively, while for the tidal Love number we find the minimum and maximum value for the whole dataset, $k^{\ast}_{2,\text{min}}$ and $k^{\ast}_{2,\text{max}}$, and we normalize each $k^{\ast}_{2,i}$ by using $(k^{\ast}_{2,i}+|k^{\ast}_{2,\text{min}}|)/(k^{\ast}_{2,\text{max}}+|k^{\ast}_{2,\text{min}}|)$. The renormalization process is an important step to render the learning process faster.
@@
III.2 Deterministic deep network
Here we construct a “standard” (deterministic) deep network, where the weights and biases of the model are computed as point estimates during the training procedure. Our model needs to be able to predict $7\times(\rho,c_{s})$ points to describe the speed-of-sound parametrization and the value of the vacuum energy shift. Hence, the output data of the neural network consist of 15 parameters; 7 values of mass density, 7 values of speed of sound, and 1 value for the vacuum energy shift. However, as we discussed in Sec. II, the EOSs we considered are generated with two different choices of low density EOS, namely AP4 and SLy, thus our neural network should also distinguish between these two classes. To achieve this goal we create two distinct models, a classification network which predicts the low-density EOS, and a regression model which predicts the remaining 15 parameters.
Our input is the data described in Sec. III.1. We first test both the classification and the regression networks for input data composed of $M$-$R$ data only. We do so in order to test our approach on a simpler problem and compare our results against those of Ref. [26]. We then tackle the full problem with a neural network that takes as input $M$-$R$-$k_{2}$ data.

### v2

@@ 1 Introduction
Recently, machine learning techniques have also been used to predict EOS’ features from observations [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Roughly, two different approaches have been followed in the literature: reconstructing the EOS from either stellar masses, stellar radii and tidal deformabilities [21, 22, 24, 25, 26, 28, 29, 30, 31, 33, 35, 37, 39, 40, 41, 43, 44, 45, 46], or focusing directly on the nuclear matter saturation properties [23, 27, 32, 34, 36, 38, 41, 42]. Our work falls into the first category. In particular, we construct a pipeline for supervised deep neural networks which allow us to predict the EOS from mock observational data composed of masses, radii and tidal deformabilities (or equivalently, tidal Love number). We employ both conventional (deterministic) and Bayesian (probabilistic) architectures. The latter utilises variational layers, where the weights and biases are drawn from a learnable posterior distribution, whereas in the deterministic network the weights and biases are point estimates. To the best of our knowledge, the employment of variational layers to account for epistemic uncertainty in the dataset is a novelty. While Bayesian neural networks have previously been used in the literature, our approach is inherently different. In particular, in Refs. [41, 42], the authors employed a Bayesian approach to minimise a log-likelihood related to neutron star observables, using a fundamentally different EOS parametrization to ours. What is more, Ref. [45] studied inferences of the EOS directly from telescope spectra employing the method of normalising flow for the reconstruction of the probability density functional.
We work in the context of a new prescriptions for modelling EOSs, following Ref. [47]. We assume that neutron stars are described by an EOS composed of three regions. At low-density we employ a realistic tabulated EOS (in this work we choose either AP4 or SLy), while at high-density we introduce a QCD phase by means of a speed-of-sound parametrization [48]. Most importantly, we also allow for a vacuum energy phase transition in the stellar core. The latter is a key novelty of our work. A vacuum energy shift, which can be interpreted as a change in the underlying value of the cosmological constant, is to be expected in the core of the star, where the pressure can be sufficiently high to trigger the QCD phase transition. This mechanism has been proposed in Ref. [49, 50, 47] to study neutron stars as novel probes for the cosmological constant problem. While it is challenging to test the evolution of the vacuum energy in the Universe, one can focus on scenarios where the structure of the vacuum is rearranged with sizable shift in the vacuum energy, as in the core of neutron stars. Hence, an imprint in the stellar properties can shed some light into the gravitational effect of such transition.
@@
The paper is structured as follows: in Section 2 we describe the modelling of the neutron star interior and the numerical pipeline we use to solve the equations which leads to our training and testing datasets. In Section 3 we describe our deep-learning pipeline, which we split into an analysis based on a deterministic and a Bayesian deep network respectively. In Section 4, we apply our network to real observational data. We summarise in Section 5. The numerical data and codes of our analysis are available online at Ref. [51].
Figure 1: A sketch showing an overview of the deep networks we use, i.e. a “standard” deterministic and a probabilistic network respectively. The input data to our network are simulated values for the mass, radius and tidal number of the given neutron star, computed according to the procedure explained in the paper, whereas the output is predominantly a prediction of the speed of sound. Our probabilistic deep network accounts for the epistemic uncertainty in the dataset, by using learnable probability functions for the weights and biases of the network, as explained in Sec. 3.3.
2 Neutron star modelling
@@
$\begin{split}m^{\prime}(r)&=4\pi r^{2}\epsilon(r)\ ,\\ p^{\prime}(r)&=-\frac{p(r)+\epsilon(r)}{r[r-2m(r)]}G[m(r)+4\pi r^{3}p(r)],\\ \nu^{\prime}(r)&=-\frac{2p^{\prime}(r)}{p(r)+\epsilon(r)}\ ,\end{split}$ (2.2)
@@
$Q_{ij}=-\lambda\mathcal{E}_{ij}=-\frac{2}{3}k_{2}R^{5}\mathcal{E}_{ij}\ ,$ (2.3)
where in the last step we have introduced the dimensionless $l=2$ Love number, $k_{2}$ [54, 55, 56, 57, 58]. The latter can be computed by considering a linear perturbation to the background solution (2.1),
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+h_{\mu\nu}(r,\theta,\phi)\ ,$ (2.4)
@@
$h_{\mu\nu}=e^{\nu(r)}H(r)Y_{20}(\theta,\phi)dt^{2}+e^{\mu(r)}H(r)Y_{20}(\theta,\phi)dr^{2}+r^{2}K(r)Y_{20}(\theta,\phi)d\Omega^{2}\ ,$ (2.5)
$T_{\mu\nu}=T_{\mu\nu}^{(0)}+\delta T_{\mu\nu},$ (2.6)
@@
$H^{\prime\prime}(r)-\frac{2}{r}e^{\mu}\big{[}2\pi Gr^{2}(\epsilon-p)-1\big{]}H^{\prime}(r)-2e^{\mu}\bigg{\{}\frac{3}{r^{2}}-2\pi G[5\epsilon+9p+(\epsilon+p)f]\\ +2G^{2}e^{\mu}\left(\frac{m(r)}{r^{2}}+4\pi rp\right)^{2}\bigg{\}}H(r)=0\ ,$ (2.7)
where $f=d\epsilon/dp$. Introducing $\beta(r)\equiv dH/dr$, we solve Eq. (2.7) integrating from the stellar center up to the surface with boundary conditions. In the inner boundary we impose
$H(r_{0})\sim a_{0}r_{0}^{2}\quad\ ,\quad\beta(r_{0})=2a_{0}r_{0}\ ,$ (2.8)
@@
$H(r\geq R)=c_{1}{\cal P}_{22}\left(r/M-1\right)+c_{2}{\cal Q}_{22}\left(r/M-1\right)\ ,$ (2.9)
@@
$-\frac{1+g_{tt}}{2}=-\frac{m}{r}-\frac{3}{2}\frac{Q_{ij}}{r^{3}}n^{i}n^{j}+...+\frac{\mathcal{E}_{ij}}{2}r^{2}n^{i}n^{j}+...,$ (2.10)
@@
$k_{2}=\frac{8C^{5}}{5}(1-2C)^{2}(2+2C(y-1)-1)\bigg{\{}2C\big{[}6-3y+3C(5y-8)\big{]}\\ +4C^{3}\big{[}13-11y+C(3y-2)+2C^{2}(1+y)\big{]}\\ +3(1-2C)^{2}\big{[}2-y+2C(y-1)\big{]}\text{ln}(1-2C)\bigg{\}}^{-1},$ (2.11)
@@
In summary, given a choice for the EOS, solving the system in Eq. (2.2) together with Eq. (2.7) allows us to extract the total stellar radius $R$, mass $M$ and dimensionless $l=2$ Love number $k_{2}$ respectively. Schematically,
$(\text{EOS})\mapsto(M,\,R,\,k_{2}).$ (2.12)
@@
A high-density regime for $\rho>\rho_{\text{tr}}$, where the nuclear matter experiences a QCD phase transition. We employ a phenomenological EOS using an agnostic speed-of-sound model [66, 48]. In particular, we construct a linear approximation of the speed of sound $c_{s}^{2}\equiv\frac{\partial p(\epsilon)}{\partial\epsilon}$, randomly sampling seven reference points $(\rho,c_{s})$ within $\rho\in(\rho_{\text{tr}},12\,\rho_{0})$ and $c_{s}\in(0,1)$, and connecting them by linear segments. Note that for the seventh point, we always assume the mass density to be fixed, i.e. $\rho_{7}=12\,\rho_{0}$, while we allow the speed of sound to acquire random values. Even though such large value for $\rho_{7}$ is unlikely to be reached by neutron stars on the stable branch, we chose to include it in our dataset to be in line with the initial assumptions of Ref. [66]. We assume causality is preserved, thus we discard models with $c_{s}>1$ to avoid unphysical EOSs.
@@
a vacuum energy phase transition in the inner stellar core, for values of the pressure larger than $p_{c}=(200\,\text{MeV})^{4}$. Indeed, deep within the stellar core, the local pressure can be sufficiently high to trigger the QCD phase transition. This would generate a shift in the underlying value of the vacuum energy. One can interpret this shift in energy as a new effective cosmological constant term $\Lambda$ contributing to the pressure, energy density and mass density of the the new exotic phase. Prior to the vacuum energy transition, the gluon condensate is thought to contribute negatively to the overall vacuum energy by an amount $-0.0034~{}(\text{GeV})^{4}$ [67]. This is the standard value usually quoted in the literature, despite some mild controversy [68, 69]. Regardless of the precise details, it is reasonable to expect the vacuum energy to increase by $(\mathcal{O}(100)\text{MeV})^{4}$ as we transition towards the inner core of the neutron star [47]. Nevertheless, we allow for both positive and negative values of $\Lambda$, to consider the possibility of an unknown particle physics mechanism screening some or all of the underlying change in vacuum energy. We follow Ref. [47], and we choose 10 values of $\Lambda$ in the range $\big{(}-(194\,\text{MeV})^{4},(194\,\text{MeV})^{4}\big{)}$, including the case where no phase transition develops in the core, i.e. $\Lambda=0$.
3 Deep learning pipeline
@@
In this section, we describe the concrete setup of our deep learning network. Our scope is to constrain the neutron star EOS from observable properties, such as the mass, the radius, and the tidal deformability of the star, or equivalently its $l=2$ dimensionless tidal Love number. Doing so corresponds to solving the “inverse” problem we delineated in Sec. 2 and described by Eq. (2.12), that is, we aim at computing
$(M,\,R,\,k_{2})\mapsto(\text{EOS}).$ (3.1)
@@
While it is known that such inverted mapping exists [70], one cannot practically retrieve it. First of all, we cannot form a continuous curve from the $M$-$R$ observational data or the $k_{2}$-$M$ ones. Furthermore, observational data have uncertainties and are described by distributions, hence one is not able to reconstruct the genuine $M$-$R$ and $k_{2}$-$M$ relations (see Ref. [26] for a further discussion on the issue). Lastly, some initial information about the EOS at low density should be known beforehand. We tackle this challenge with an alternative approach, similar to Ref. [26]. In particular, instead of seeking the unique inverse mapping, i.e. Eq. (3.1), we infer the most likely EOS from observations data including their uncertainties through a regression analysis with deep-learning techniques. We do this at two steps. We first consider a “standard” (deterministic) deep network. In this case, the weights and biases of the network at training are point estimates, which implies that once the network is trained, a given input will always output the same prediction, modulo the inherit stochasticity in deep networks. Our second approach will be a Bayesian deep network, able to capture the epistemic uncertainty in the data. In this approach, the weights and biases at each node are drawn from learnable probability density functions, and the output of the network will be a statistical realisation of the given input. Before we dive into the description of the deep architecture and their results, we first discuss the construction of the training and validation dataset.
3.1 Training and validation dataset
@@
In order to train and test both the deterministic and the Bayesian neural networks, we first generate mock observational data as input data. We first solve the TOV equations (2.2) for randomly generated EOSs (see Sec. 2), obtaining for each EOS a $M$-$R$ and a $k_{2}$-$M$ curve respectively. Note that, in order to generate data in agreement with the heaviest neutron star observed so far (PSR J0952-0607 [71]), which has a mass $M_{\text{max}}=(2.35\pm 0.17)M_{\odot}$, we discard all EOSs which predict a maximum mass not included in this interval. For each EOS, we then randomly sample $N$ mock observation points $(M_{i},R_{i},k_{2,i})$, assuming a uniform distribution of neutron star masses in the interval $(M_{\text{min}},M_{\text{max}})$, where $M_{\text{min}}$ is the minimum mass produced by the given EOS  [^2] The choice of this interval allow us to discard unstable branches of the $M$-$R$ curve, while associating each $M$ to a single value of $R$ and $k_{2}$.. We initially chose $N=15$, following Ref. [26], but we also took into consideration the case of a larger number of mock data for a given EOS, i.e. we consider $N=20$ and $N=30$. As we will discuss in Sec. 3.2, allowing for a larger set of observation points for each $M$-$R$ and $k_{2}$-$M$ curve increases the accuracy of our network.
@@
As we discussed in Sec. 2, we generate $2\times 10^{4}$ EOSs, of which only $1491$ survive the maximum mass test. For each set of points $N\times(M^{\ast}_{i},R^{\ast}_{i},k^{\ast}_{2,i})$, we create $n$ different realisations through noise injection, leading to $N\times n\times 3$ points in total for each EOS. All values are normalized appropriately by employing $M^{\ast}_{i}/M_{\text{norm}}$, and $R^{\ast}_{i}/R_{\text{norm}}$ with $M_{\text{norm}}=3\,M_{\odot}$ and $R_{\text{norm}}=20\,\text{km}$ for masses and radii respectively, while for the tidal Love number we find the minimum and maximum value for the whole dataset, $k^{\ast}_{2,\text{min}}$ and $k^{\ast}_{2,\text{max}}$, and we normalize each $k^{\ast}_{2,i}$ by using $(k^{\ast}_{2,i}+|k^{\ast}_{2,\text{min}}|)/(k^{\ast}_{2,\text{max}}+|k^{\ast}_{2,\text{min}}|)$. The renormalization process is an important step to render the learning process faster.
@@
3.2 Deterministic deep network
Here we construct a “standard” (deterministic) deep network, where the weights and biases of the model are computed as point estimates during the training procedure. Our model needs to be able to predict $7\times(\rho,c_{s})$ points to describe the speed-of-sound parametrization and the value of the vacuum energy shift. Hence, the output data of the neural network consist of 15 parameters; 7 values of mass density, 7 values of speed of sound, and 1 value for the vacuum energy shift. However, as we discussed in Sec. 2, the EOSs we considered are generated with two different choices of low density EOS, namely AP4 and SLy, thus our neural network should also distinguish between these two classes. To achieve this goal we create two distinct models, a classification network which predicts the low-density EOS, and a regression model which predicts the remaining 15 parameters.
Our input is the data described in Sec. 3.1. We first test both the classification and the regression networks for input data composed of $M$-$R$ data only. We do so in order to test our approach on a simpler problem and compare our results against those of Ref. [26]. We then tackle the full problem with a neural network that takes as input $M$-$R$-$k_{2}$ data.